# Distributed Maximum Flow in Planar Graphs

### v3

@@ 2.2 SSSP in the Dual via Distance Labels
There is a randomized distributed $\tilde{O}(D^{2})$-round algorithm that computes $\tilde{O}(D)$-bit distance labels for $G^{*}$, or reports that $G^{*}$ contains a negative cycle w.h.p., upon termination, each vertex of $G$ that lies on a face $f$, knows the label of the node $f$ in $G^{*}$.
@@
As known from the 70’s [27, 29], the family of planar graphs admits cycle separators (a cycle whose removal disconnects the graph) of small size $O(\sqrt{n})$ or $O(D)$. Since $\Omega(D)$ is a lower bound for all the problems we discuss in the distributed setting, we will focus on separators of size $\tilde{O}(D)$. The centralized divide-and-conquer approach repetitively removes the separator vertices $S_{\!{}_{G}}$ from the graph $G$ and recurses on the two remaining subgraphs that are a constant factor smaller, constituting a hierarchical decomposition of the graph with $O(\log n)$ levels. The labeling scheme is defined recursively, where the label of a vertex $v$ in $G$ stores distances between $v$ and $S_{\!{}_{G}}$ and recursively the label of $v$ in the subgraph that contains $v$ (either the interior or the exterior of $S_{\!{}_{G}}$). Due to $S_{\!{}_{G}}$ being a cycle, any shortest $u$-to-$v$ path $P$ for any two vertices $u,v$ in $G$ either crosses $S_{\!{}_{G}}$ (and the $u$-to-$v$ distance is known by the distances between $u,v$ and $S_{\!{}_{G}}$ stored in their labels), or $P$ is enclosed in the interior or the exterior of $S_{\!{}_{G}}$ (and the $u$-to-$v$ distance is decoded from the recursive labels of $u$ and $v$ in the subgraph that contains them). Moreover, since the separator is of size $\tilde{O}(D)$ and since there are $O(\log n)$ levels in the decomposition, the labels are of size $\tilde{O}(D)$.
@@
Figure 1: Two levels of the recursive separator decomposition of $G$ and their correspondent levels in the desired dual decomposition of $G^{*}$. The black edges are the separating cycle $S_{\!{}_{G}}$ in $G$, the same edges map to a cut in $G^{*}$. For simplicity, the dual node corresponding to the face $f_{\infty}$ of $G$ is not illustrated in $G^{*}$ (its incident edges are the ones missing an endpoint).
Our approach: recurse on primal, solve on dual. To overcome this, we suggest a different approach. We use the primal BDD, and infer from it a decomposition of the dual graph. To get an intuition, note that by the cycle-cut duality, a cycle in $G$ constitute a cut in $G^{*}$. This means that the separator of $G$ can be used as a separator of $G^{*}$, in the sense that removing the cut edges (or their endpoints) disconnects $G^{*}$. I.e, a hierarchical decomposition of $G$ can (conceptually) be thought of as a hierarchical decomposition of $G^{*}$. See Figure 1.
Then, we would hope that using [26]’s algorithm for such a hierarchical decomposition for planar graphs would allow us to apply the same labeling scheme on $G^{*}$, where we consider the dual endpoints of $S_{\!{}_{G}}$ edges as our separator in $G^{*}$. However, taking a dual lens on the primal decomposition introduces several challenges that arise when one needs to define the dual subgraphs from the given primal subgraphs. This primal to dual translation is rather non-trivial due to critical gaps that arise when one needs to maintain information w.r.t. faces of $G$, rather than vertices of $G$. To add insult to injury, the nice structural property of the separator being a cycle is simply not true any more. This is because the separator of [17] that the BDD of [26] uses is composed of two paths plus an additional edge that exists only when the graph is two-connected. Otherwise, that edge is not in $G$ and cannot be used for communication. This virtual edge, if added, would split a face into two parts, introducing several challenges, including:
@@
Challenge II: face-parts. In the primal graph, a vertex is an atomic unit, which keeps its identity throughout the computation. The situation in the dual graph is considerably more involved. Consider a primal graph $G$ with a large face $f$ containing $\Theta(n)$ edges. Throughout the recursive decomposition, the vertices of the face $f$ are split among multiple faces, denoted hereafter as face-parts, and eventually $f$ is shattered among possibly a linear number of subgraphs. This means that a node in a dual subgraph does no longer correspond to a face $f$ of the primal subgraph, but rather to a subset of edges of $f$. This creates a challenge in the divide-and-conquer computation, where one needs to assemble fragments of information from multiple subgraphs.
@@
Figure 3: Three examples of a bag $X$ and its child bags $X_{i}$ (of distinct colors). The bold black cycle is the separator $S_{\!{}_{X}}$. The virtual edge $e_{\!{}_{X}}$ is dashed. The green child bag $X_{1}$ is the interior of $S_{\!{}_{X}}$. When $e_{\!{}_{X}}\in E(G)$ (left image), we show that no face is partitioned between the child bags $X_{i}$. When $e_{\!{}_{X}}\notin E(G)$ (middle and right images), edges of the critical face $f$ are red. Notice, the face $f$ (middle) is partitioned into two child bags ($X_{1}$ and $X_{4}$) and the face-part $f$ (right) is partitioned into three child bags ($X_{1},X_{4},X_{5}$).
@@
To prove Lemma 2.3, we consider the separator $S_{\!{}_{X}}$ of $X$ computed in the construction of the BDD by [26]. $S_{\!{}_{X}}$ constitutes a separating cycle in $X$ containing at most one virtual edge $e_{\!{}_{X}}$ that might not be in $E(G)$. The interior of $S_{\!{}_{X}}$ defines one child bag $X_{1}$ of $X$, and its exterior may define as many as $\tilde{O}(D)$ child bags $X_{2},X_{3},\ldots$ (see Figure 3). We prove that the face containing the endpoints of $e_{\!{}_{X}}$ is the only face that might get partitioned in the bag $X$. First, we show that if $e_{\!{}_{X}}\in E(G)$ then no face is partitioned, and all faces are entirely contained in one of the child bags. If $e_{\!{}_{X}}\notin E(G)$, then we show that the critical face is the only face of $G$ that can be partitioned in $X$. We remark that face-parts can also be further partitioned, as shown in Figure 3.
Dual bags and separators. Since we are interested in faces of $G$, we trace them down the decomposition and take face-parts into consideration when defining the dual bags. That is, both the faces and the face-parts that belong to a bag $X$ have corresponding dual nodes in $X^{*}$. We connect two nodes of $X^{*}$ by a dual edge only if their corresponding faces share a primal (non-virtual) edge in $X$. If $X=G$, then there are no face-parts and $X^{*}$ is the standard dual graph $G^{*}$. For full details, the reader is referred to Section 5.1.
@@
The labeling scheme. Our next goal is to use the decomposition in order to compute distances in the dual graph. More concretely, we (distributively) compute, for every node in every bag $X^{*}$, a short label of size $\tilde{O}(D)$, such that given the labels alone of any two nodes in $X^{*}$ we can deduce their distance. The labeling scheme is a refinement of the intuition given earlier. For full details, the reader is referred to Section 5.2.
Recall, the set $F_{\!{}_{X}}$ plays the role of a node separator of $X^{*}$ (Lemma 2.4). A node $g\notin F_{\!{}_{X}}$ of $X^{*}$ corresponds to a real face of $G$ (i.e. not a face-part) contained in $X$. The distance label $\mathsf{Label}_{X^{*}}(g)$ of $g\notin F_{\!{}_{X}}$ is defined recursively. If $X^{*}$ is a leaf in the decomposition then $\mathsf{Label}_{X^{*}}(g)$ stores ID($g$) and the distances between $g$ and all other nodes $h\in X^{*}$. Otherwise, the label consists of the ID of $X$, the distances in $X^{*}$ between $g$ and all nodes of $F_{\!{}_{X}}$, and (recursively) the label of $g$ in the child bag of $X^{*}$ that entirely contains $g$. In case $g\in F_{\!{}_{X}}$, we simply store the distances between $g$ and all other nodes in $F_{\!{}_{X}}$. The labels are of size $\tilde{O}(D)$-bits since $|F_{\!{}_{X}}|=\tilde{O}(D)$ and since the BDD is of a logarithmic height.
[…]
@@ 2.3 Flow Assignments and Cuts
Maximum $st$-flow. It was shown in the 90’s by Miller and Naor [30], that an exact $st$-flow algorithm on a directed planar graph $G$ with positive edge-capacities can be obtained by $\log\lambda$ applications of a SSSP algorithm on the dual graph $G^{*}$ with positive and negative edge-weighs (each application with possibly different edge weights). We use this result as a black-box. At the end, we obtain the flow value and the flow assignment to the edges of $G$. This works in $\tilde{O}(D^{2})$ rounds, since $\lambda$ (the value of the maximum $st$-flow) is polynomial in $n$ and since our dual SSSP algorithm terminates within $\tilde{O}(D^{2})$ rounds.
Approximate maximum $st$-flow. For the case of undirected planar graphs where $s$ and $t$ lie on the same face with respect to the given embedding, we further obtain an improved round complexity at the cost of solving the problem approximately. Namely, it was shown in the 80’s by Hassin [20] that in this setting it suffices to use an SSSP algorithm on the dual graph after augmenting it with a single edge, dual to the edge $(s,t)$, that is limited to positive edge-weights and to undirected graphs. Since $s$ and $t$ are on the same face it is possible to add such an edge while preserving planarity. Note that adding this edge splits a node in the dual graph. We show that we can still simulate efficently minor-aggregation algorithms on the augmented dual graph. To compute SSSP on that graph we use an SSSP algorithm by [42] that works in the minor-aggregation model. In particular, it computes a $(1+\epsilon)$-approximate SSSP tree within $D\cdot n^{o(1)}$ rounds, which allows us to find a $(1-\epsilon)$-approximate flow value in this time.
@@
Minimum $st$-cut. By the celebrated min-max theorem of Ford and Fulkerson [7], the exact (or approximate) value of the max $st$-flow is equivalent to that of the min $st$-cut. We show that our algorithms can also be extended to compute a corresponding bisection and mark the cut edges of the min $st$-cut. In the exact case, we use classic textbook methods (e.g, residual graphs), and since this flow algorithm works for directed planar graphs, we obtain a directed minimum $st$-cut. In the approximate case, we have to work harder and exploit the minor-aggregation model on the dual graph one last time, in which we apply a special argument on the cycle-cut duality that was given in the 80’s by Reif [38].
Directed global minimum cut. We exploit our techniques to find also the directed global minimum cut. The general idea is simple, we use the cycle-cut duality: finding the (directed) minimum cut in $G$ amounts to finding the minimum weight (directed) simple cycle $C$ in $G^{*}$. Thus, given an extended BDD of $G$ and having distance labels already computed in $\tilde{O}(D^{2})$ rounds for all dual bags in the BDD, we have that any shortest cycle of $G^{*}$ either intersects a dual separator node or is entirely contained in one of $G^{*}$’s child bags. Thus, in order to find $C$, we take the minimum over all: (i) directed cycles that intersect separator nodes of $G^{*}$ (deduced by pairwise decoding labels of the separator nodes), and (ii) minimum directed cycles in $G^{*}$’s child bags (found recursively). However, we note that the duality does not hold in general directed planar graphs, which complicates the above approach. I.e., in order for the duality to hold, we have to work with a graph where for each edge, its reversal dart (an edge in the opposite direction) is present, so we add those darts. Then, we should ensure that the cycles we consider are simple darts cycles (i.e. a cycle where no edge and its reversal dart appear together), which we do by investigating two options related to the dual separator. Finding the cut bisection is done simply by detecting connected components (in the undirected sense) after the removal of the cut edges and their reversal darts.
@@ 2.4 Organization, Summary and Open Questions
The rest of the paper is organized as follows. Section 3 discusses preliminaries. Section 4 gives an overview of our minor-aggregation simulation on the dual graph and its applications for the weighted girth, where full details and proofs appear in Appendix B. Our labeling scheme for the dual graph appears in Section 5, and our applications for max $st$-flow, min $st$-cut appear in Section 6, while the directed global min-cut appears in Section 7. See also Figure 6.
Dual Minor-Aggregation (Section 4, Appendix B)Undirected Weighted Girth (Theorem 1.7)Approximate Dual SSSPExtended BDD (Section 5.1)Dual Distance Labeling (Section 5.2)Exact Dual SSSP(Section 5.3)Primal Max $st$-Flow(Section 6.1)Primal Directed Global Min-Cut(Section 7)Primal Min $st$-Cut(Section 6.2)
@@
This closes the problem of minor-aggregation on the dual and expands the algorithmic design toolkit to include all the very useful efficient minor-aggregation procedures (e.g. subtree sums, Heavy-Light decompositions [18], etc) and even complicated algorithms (e.g. min-cut [18], approximate undirected shortest path [41]) for the dual.

### v4

@@ 2.2 SSSP in the Dual via Distance Labels
There is a randomized distributed $\tilde{O}(D^{2})$-round algorithm that computes $\tilde{O}(D)$-bit distance labels for $G^{*}$, or reports that $G^{*}$ contains a negative cycle w.h.p.. Upon termination, each vertex of $G$ that lies on a face $f$, knows the label of the node $f$ in $G^{*}$.
@@
As known from the 70’s [28, 30], the family of planar graphs admits cycle separators (a cycle whose removal disconnects the graph) of small size $O(\sqrt{n})$ or $O(D)$. Since $\Omega(D)$ is a lower bound for all the problems we discuss in the distributed setting, we will focus on separators of size $\tilde{O}(D)$. The centralized divide-and-conquer approach repetitively removes the separator vertices $S_{\!{}_{G}}$ from the graph $G$ and recurses on the two remaining subgraphs that are a constant factor smaller, constituting a hierarchical decomposition of the graph with $O(\log n)$ levels. The labeling scheme is defined recursively, where the label of a vertex $v$ in $G$ stores distances between $v$ and $S_{\!{}_{G}}$ and recursively the label of $v$ in the subgraph that contains $v$ (either the interior or the exterior of $S_{\!{}_{G}}$). Due to $S_{\!{}_{G}}$ being a cycle, any shortest $u$-to-$v$ path $P$ for any two vertices $u,v$ in $G$ either crosses $S_{\!{}_{G}}$ (and the $u$-to-$v$ distance is known by the distances between $u,v$ and $S_{\!{}_{G}}$ stored in their labels), or $P$ is enclosed in the interior or the exterior of $S_{\!{}_{G}}$ (and the $u$-to-$v$ distance is decoded from the recursive labels of $u$ and $v$ in the subgraph that contains both of them). Moreover, since the separator is of size $\tilde{O}(D)$ and since there are $O(\log n)$ levels in the decomposition, the labels are of size $\tilde{O}(D)$.
@@
Figure 2: Two levels of the recursive separator decomposition of $G$ and their correspondent levels in the desired dual decomposition of $G^{*}$. The black edges demonstrate the separating cycle $S_{\!{}_{G}}$ in $G$, the same edges map to a cut in $G^{*}$. For simplicity, the dual node corresponding to the face $f_{\infty}$ of $G$ is not illustrated in $G^{*}$ (its incident edges are the ones missing an endpoint).
Our approach: recurse on primal, solve on dual. To overcome this, we suggest a different approach. We use the primal BDD, and infer from it a decomposition of the dual graph. To get an intuition, note that by the cycle-cut duality, a cycle in $G$ constitutes a cut in $G^{*}$. This means that the separator of $G$ can be used as a separator of $G^{*}$, in the sense that removing the cut edges (or their endpoints) disconnects $G^{*}$. I.e, a hierarchical decomposition of $G$ can (conceptually) be thought of as a hierarchical decomposition of $G^{*}$. See Figure 2.
Then, we would hope that using [27]’s algorithm for such a hierarchical decomposition for planar graphs would allow us to apply the same labeling scheme on $G^{*}$, where we consider the dual endpoints of $S_{\!{}_{G}}$ edges as our separator in $G^{*}$. However, taking a dual lens on the primal decomposition introduces several challenges that arise when one needs to define the dual subgraphs from the given primal subgraphs. This primal to dual translation is rather non-trivial due to critical gaps that arise when one needs to maintain information w.r.t. faces of $G$, rather than vertices of $G$. To add insult to injury, the nice structural property of the (primal) separator being a cycle of $G$ is simply not true any more. This is because the separator of [17] that the BDD of [27] uses is composed of two paths plus an additional edge that exists only when the graph is two-vertex connected. Otherwise, that edge is not in $G$ and cannot be used for communication. This virtual edge, if added, would split a face into two parts, introducing several challenges, including:
@@
Challenge II: face-parts. In the primal graph, a vertex is an atomic unit, which keeps its identity throughout the computation. The situation in the dual graph is considerably more involved. Consider a primal graph $G$ with a large face $f$ containing $\Theta(n)$ edges. Throughout the recursive decomposition, the vertices of the face $f$ are split among multiple faces, denoted hereafter as face-parts, and eventually $f$ is shattered among possibly a linear number of subgraphs. This means that a node in a dual subgraph no longer corresponds to a face $f$ of the primal subgraph, but rather to a subset of edges of $f$. This creates a challenge in the divide-and-conquer computation, where one needs to assemble fragments of information from multiple subgraphs.
@@
Figure 4: Three examples of a bag $X$ and its child bags $X_{i}$ (of distinct colors). The bold black cycle is the separator $S_{\!{}_{X}}$. The virtual edge $e_{\!{}_{X}}$ is dashed. The green child bag $X_{1}$ is the interior of $S_{\!{}_{X}}$. When $e_{\!{}_{X}}\in E(G)$ (left image), we show that no face is partitioned between the child bags $X_{i}$. When $e_{\!{}_{X}}\notin E(G)$ (middle and right images), edges of the critical face $f$ are red. Notice, the face $f$ (middle) is partitioned into two child bags ($X_{1}$ and $X_{4}$), and the face-part $f$ (right) is further partitioned into three child bags ($X_{1},X_{4},X_{5}$); I.e., in the right image, $f$ was a face in a previous levels that got partitioned into face-parts, in particular, each dashed edge $e_{\!{}_{H}}\neq e_{\!{}_{X}}$ that appears in the empty white subgraph is an edge of a separator $S_{\!{}_{H}}$ from an ancestor bag $H$ of $X$, such that $e_{\!{}_{H}}$ had partitioned $f$ into parts.
@@
To prove Lemma 2.3, we consider the separator $S_{\!{}_{X}}$ of $X$ computed in the construction of the BDD by [27]. $S_{\!{}_{X}}$ constitutes a separating cycle in $X$ containing at most one virtual edge $e_{\!{}_{X}}$ that might not be in $E(G)$. The interior of $S_{\!{}_{X}}$ defines one child bag $X_{1}$ of $X$, and its exterior may define as many as $\tilde{O}(D)$ child bags $X_{2},X_{3},\ldots$ (see Figure 4). We prove that the face containing the endpoints of $e_{\!{}_{X}}$ is the only face that might get partitioned in the bag $X$. First, we show that if $e_{\!{}_{X}}\in E(G)$ then no face is partitioned, and every face is entirely contained in one of the child bags. If $e_{\!{}_{X}}\notin E(G)$, then we show that the critical face is the only face of $G$ that can be partitioned in $X$. We remark that face-parts can also be further partitioned, as shown in Figure 4.
Dual bags and separators. Since we are interested in faces of $G$, we trace them down the decomposition and take face-parts into consideration when defining the dual bags. That is, both the faces and the face-parts that belong to a bag $X$ have corresponding dual nodes in $X^{*}$. We connect two nodes of $X^{*}$ by a dual edge only if their corresponding faces share a primal (non-virtual) edge in $X$. If $X=G$, then there are no face-parts, and $X^{*}$ is the standard dual graph $G^{*}$. For full details, the reader is referred to Section 5.1.
@@
The labeling scheme. Our next goal is to use the decomposition in order to compute distances in the dual graph. More concretely, we (distributively) compute, for every node in every bag $X^{*}$, a short label of size $\tilde{O}(D)$, such that given only the labels of any two nodes in $X^{*}$ we can deduce their distance. The labeling scheme is a refinement of the intuition given earlier. For full details, the reader is referred to Section 5.2.
Recall, the set $F_{\!{}_{X}}$ plays the role of a node separator of $X^{*}$ (Lemma 2.4). A node $g\notin F_{\!{}_{X}}$ of $X^{*}$ corresponds to a real face of $G$ (i.e. not a face-part) contained in $X$. The distance label $\mathsf{Label}_{X^{*}}(g)$ of $g\notin F_{\!{}_{X}}$ is defined recursively. If $X^{*}$ is a leaf in the decomposition then $\mathsf{Label}_{X^{*}}(g)$ stores ID($g$) and the distances between $g$ and all other nodes $h\in X^{*}$. Otherwise, the label consists of the ID of XX [^6] In the BDD of [27], each bag is assigned a $O(\log n$)-bit identifier., the distances in $X^{*}$ between $g$ and all nodes of $F_{\!{}_{X}}$, and (recursively) the label of $g$ in the child bag of $X^{*}$ that entirely contains $g$. In case $g\in F_{\!{}_{X}}$, we simply store the distances between $g$ and all other nodes in $F_{\!{}_{X}}$. The labels are of size $\tilde{O}(D)$-bits since $|F_{\!{}_{X}}|=\tilde{O}(D)$ and since the BDD is of a logarithmic height.
The correctness follows from Lemma 2.4. I.e, a $g$-to-$h$ shortest path $P$ in $X^{*}$ is either: (1) Entirely contained in $X^{*}_{i}$, and $\mathsf{dist}_{X^{*}}(g,h)$ can be deduced from $\mathsf{Label}_{X^{*}_{i}}(g)$ and $\mathsf{Label}_{X^{*}_{i}}(h)$, or (2) $P$ intersects $F_{\!{}_{X}}$ with an edge, then $\mathsf{dist}_{X^{*}}(g,h)=\min_{f\in F_{\!{}_{X}}}\{\mathsf{dist}_{X^{*}}(g,f)+\mathsf{dist}_{X^{*}}(f,h)\}$. If one of $g$ or $h$ is a node of $F_{\!{}_{X}}$ then distances are retrieved instantly, as each node in $X^{*}$ stores its distances to $F_{\!{}_{X}}$.
@@ 2.3 Flow Assignments and Cuts
Maximum $st$-flow. It was shown in the 90’s by Miller and Naor [31], building on a result of Venkatesan [46], that an exact $st$-flow algorithm on a directed planar graph $G$ with positive edge-capacities can be obtained by $\log\lambda$ applications of a SSSP algorithm on the dual graph $G^{*}$ with positive and negative edge-weighs (each application with possibly different edge weights). We use this result as a black-box. At the end, we obtain the flow value and the flow assignment to the edges of $G$. This works in $\tilde{O}(D^{2})$ rounds since $\lambda$ (the value of the maximum $st$-flow) is polynomial in $n$, and since our dual SSSP algorithm terminates within $\tilde{O}(D^{2})$ rounds.
Approximate maximum $st$-flow. For the case of undirected planar graphs where $s$ and $t$ lie on the same face with respect to the given embedding, we further obtain an improved round complexity at the cost of solving the problem approximately. Namely, it was shown in the 80’s by Hassin [20] that in this setting it suffices to use an SSSP algorithm on the dual graph after augmenting it with a single edge, dual to the edge $(s,t)$, that is limited to positive edge-weights and to undirected graphs. Since $s$ and $t$ are on the same face it is possible to add such an edge while preserving planarity. Note that adding this edge splits a node in the dual graph. We show that we can still simulate efficently minor-aggregation algorithms on the augmented dual graph. To compute SSSP on that graph we use an SSSP algorithm by [43] that works in the minor-aggregation model. In particular, it computes a $(1+o(1))$-approximate SSSP tree within $D\cdot n^{o(1)}$ rounds, which allows us to find a (1o(1))(1-o(1))-approximate flow value in the same time bound.
@@
Minimum $st$-cut. By the celebrated min-max theorem of Ford and Fulkerson [7], the exact (or approximate) value of the max $st$-flow is equivalent to that of the min $st$-cut. We show that our algorithms can also be extended to compute a corresponding bisection and mark the cut edges of the min $st$-cut. In the exact case, we use classic textbook methods (e.g, residual graphs), and since this flow algorithm works for directed planar graphs, we obtain a directed minimum $st$-cut. In the approximate case, we have to work harder and exploit the minor-aggregation model on the dual graph one last time, in which, we apply a special argument on the cycle-cut duality that was given in the 80’s by Reif [39].
Directed global minimum cut. We exploit our techniques to also find the directed global minimum cut. The general idea is simple, we use the cycle-cut duality: finding the (directed) minimum cut in $G$ amounts to finding the minimum weight (directed) simple cycle $C$ in $G^{*}$. Thus, given an extended BDD of $G$ and having distance labels already computed in $\tilde{O}(D^{2})$ rounds for all dual bags in the BDD, we have that any shortest cycle of $G^{*}$ either intersects a dual separator node or is entirely contained in one of $G^{*}$’s child bags. Thus, in order to find $C$, we take the minimum over all: (i) directed cycles that intersect separator nodes of $G^{*}$ (deduced by pairwise decoding labels of the separator nodes), and (ii) minimum directed cycles in $G^{*}$’s child bags (found recursively). However, we note that the duality does not hold in general directed planar graphs, which complicates the above approach. I.e., in order for the duality to hold, we have to work with a graph where for each edge, its reversal dart (an edge in the opposite direction) is present, so we add those darts. Then, we should ensure that the cycles we consider are simple darts cycles (i.e. a cycle where no edge and its reversal dart appear together), which we do by investigating two options related to the dual separator. Finding the cut bisection is done simply by detecting connected components (in the undirected sense) after the removal of the cut edges and their reversal darts.
@@ 2.4 Organization, Summary and Open Questions
The rest of the paper is organized as follows. Section 3 discusses preliminaries. Section 4 gives the proofs for our minor-aggregation simulation on the dual graph and its applications for the weighted girth. Our labeling scheme for the dual graph appears in Section 5, and our applications for max $st$-flow, min $st$-cut appear in Section 6, while the directed global min-cut appears in Section 7. See also Figure 7.
Dual Minor-Aggregation (Section 4)Undirected Weighted Girth (Theorem 1.7)Approximate Dual SSSPExtended BDD (Section 5.1)Dual Distance Labeling (Section 5.2)Exact Dual SSSP(Section 5.3)Primal Max stst-Flow(Section 6.1)Primal Directed Global Min-Cut(Section 7)Primal Min stst-Cut(Section 6.2)
@@
This closes the problem of minor-aggregation on the dual and expands the algorithmic design toolkit to include all the very useful efficient minor-aggregation procedures (e.g. subtree sums, heavy-light decompositions [18], and more) and even complicated algorithms (e.g. min-cut [18], approximate undirected shortest path [43]) for the dual.